Two angles are complementary if the sum of their measures is 90°. Find two complementary angles such that one of the angles is 142° less than 3 times the other angle. (Round to two decimal places if necessary.)
_____ the smaller complementary angle?
---- the larger complementary angle?

Answer by ankor@dixie-net.com(22740)

Two angles are complementary if the sum of their measures is 90°.
Find two complementary angles such that one of the angles is 142° less than 3 times the other angle.
A + B = 90
and
A = 3B - 142
A - 3B = -142
:
Use elimination on the following two equations
A + B = 90
A - 3B = -142
-----------------Subtraction eliminates A, find B
0 + 4B = 232
B = 232/4
B = 58 degrees
and
90 - 58 = 32 degrees

__32___ the smaller complementary angle?
__58___ the larger complementary angle?